# Is the Lecture Engaging for Learning? Lecture Voice Sentiment Analysis for Knowledge Graph-Supported Intelligent Lecturing Assistant (ILA) System

### v1

@@ 8 Experimental Results
Table 4 presents the best evaluation results for each model type specified in Table 3. Specifically, Table 4 lists the parameters of the optimal model and the feature combinations that yielded the highest F1-measure score for the boring label. Additionally, the table includes accuracy and other metrics for the engaging label for reference. In Table 4, the metric Precision_1 refers to the metric calculated for label 1 (boring). This interpretation applies similarly to other metrics with suffixes.
@@
denseNet See Table 3 zcrate_mean 0.81 0.91 0.80 0.86 0.65 0.83 0.73
@@
Table 4: The Best Evaluation Results of Each Type of Model and Features.
@@ 9 Discussion
This paper introduces the application of knowledge graphs at a high level, with a primary focus on lecture voice sentiment analysis. However, the knowledge graph functions as the "brain" of the intelligent lecturing assistant system, offering the contextual information and reasoning capabilities necessary to act on insights derived from sentiment analysis. This information is then utilized by the system to make informed decisions regarding when and how to intervene to enhance student engagement and learning outcomes. The construction and implementation of the knowledge graph for reasoning within the lecturing assistant system remain areas for further development.
Ethical Considerations: The deployment of AI in educational settings raises important ethical issues. Future research must address concerns related to data bias, privacy, and the potential impacts of AI on existing inequalities.
Intelligent Lecturing Assistant (ILA) vs. Intelligent Tutoring System (ITS): This paper proposes an Intelligent Lecturing Assistant (ILA) system designed to enhance learning engagement and the effectiveness of lectures delivered by human instructors. In contrast, Intelligent Tutoring Systems (ITS) are primarily developed to offer personalized instruction and guidance to individual learners, often within a self-paced learning environment. While these systems serve distinct purposes, ILAs and ITSs have the potential to complement one another in improving students’ learning outcomes.
10 Conclusion

### v2

@@ VIII Experimental Results
Table IV presents the best evaluation results for each model type specified in Table III. Specifically, Table IV lists the parameters of the optimal model and the feature combinations that yielded the highest F1-measure score for the boring label. Additionally, the table includes accuracy and other metrics for the engaging label for reference. In Table IV, the metric Precision_1 refers to the metric calculated for label 1 (boring). This interpretation applies similarly to other metrics with suffixes.
@@
denseNet See Table III zcrate_mean 0.81 0.91 0.80 0.86 0.65 0.83 0.73
@@
TABLE IV: The Best Evaluation Results of Each Type of Model and Features.
@@ IX Discussion
Ethical Considerations: The analysis of lecture voice sentiment raises several important ethical issues that must be carefully considered. Firstly, the datasets could introduce biases inherent to the specific context in which the data was gathered. Such biases might limit the generalizability of the findings to different educational environments, demographics, or cultural contexts. Future research should strive to mitigate these concerns. Secondly, the collection and analysis of lecture voice data involve sensitive information that could potentially be misused if not properly safeguarded. Ensuring the privacy is paramount. Finally, the introduction of AI into educational contexts has the potential to exacerbate existing inequalities. It is crucial that AI systems are designed with inclusivity in mind, ensuring that they are fair and equitable across diverse teaching styles and populations.
X Conclusion